Started the troubleshoot from the MC1 issue. Gautam showed me how to use the fake PD/LED pair to diagnose the satellite box without involving the suspension mechanics.
This revealed that the MC1 has frequent light level glitches which are common for five sensors. This feature does not exist in the test with the MC3 satellite box. I will open and check the MC1 satellite box to find the cause of this common glitches tomorrow. MC1 is currently shutdown and undamped.
BTW, at the MC3 test, i found that J2 of the satellite box (male Dsub) has all the pins too low (or too short?). I brought the box outside and found that the housing of this connector was half broken down. The connector was reassembled and the metal parts of the housing was bent again so that the housing can hold the connector body tightly.
The MC3 satellite box was restored and connected to the cables. As I touched this box, it is still under probation.
Here is what is left to do:
Today I set up the autoburt.req file for the c1iscaux channels, and confirmed that the snapshots are getting recorded. There were a lot of channels in the old autoburt.req file which I thought were un-necessary (and several which no longer exist), so now the only channels that are burt-ed are the whitening gains and states of the AA filters. If someone feels we need more channels to be snapshot recorded, you can add them to the file.
In the old target directory, there were also various versions of a "saverestore.req" file - why do we need this in addition to an autoburt? I guess it is possible they are used by the IFOconfigure scripts to setup some whitening gains etc...
Work done today:
Testing of functionality:
Much testing remains to be done, but I defer further testing till Monday - the main functionality to be verified in the short run is the whitening gain stepping. The strain-relief of cables and general cleanup will be undertaken by Chub. Current state of affairs is in Attachment #3, leaves much to be desired in terms of cleanliness.
I will also setup the autoburt for the new machine on Monday. We will also need to add some channels to C0EDCU.ini if we want to trend them over some years (e.g. RF signal powers for monitoring ERA-5 health).
* c1lsc FE was rebooted using the usual script, and everything seems to be healthy in CDS-land again, see Attachment #4.
Installation: The following equipment were installed in 1Y3, see Attachment #1:
Removal: The following equipment was removed from 1Y3:
I judge that we are good to go ahead with an install tomorrow.
What about just use high gain feedback to MC2 below 20 Hz for the IMC lock? That would reduce the excess if this theory is correct.
I came aross an interesting suggestion by Yutaro that KAGRA's low-frequency ALS noise could be limited by the fact that the IMC comes between the point where the frequencies of the PSL and AUX lasers are sensed (i.e. the ALS beat note), and the point where we want them to be equal (i.e. the input of the arm cavity). I wanted to see if the same effect could be at play in the 40m ALS system. A first estimate suggests to me that the numbers are definitely in the ballpark. If this is true, we may benefit from lower noise ALS by picking off the PSL beam for the ALS beat note after the IMC.
Even though the KAGRA phase lock scheme is different from the 40m scheme, the algebra holds. I needed an estimate of how much the arm cavity moves, I used data from a POX lock to estimate this. The estimate is probably not very accurate (since the arm cavity length is more stable than the IMC length, and the measured ALS noise, e.g. this elog, is actually better than what this calculation would have me believe), but should be the right order of magnitude. From this crude estimate, it does look like for f<10 Hz, this effect could be significant. I assumed an IMC pole of 3.8 kHz for this calculation.
I've indicated a "target" ALS performance where the ALS noise would be less than the CARM linewidth, which would hopefully make the locking much easier. Seems like realizing this target will be touch-and-go. But if we can implement length feedforward control for the arm cavities using seismometers, the low frequency motion of the optics should go down. It would be interesting to see if the ALS noise gets better at low frequencies with length feedforward engaged.
* Some updates were made to the plot:
As it turns out, only one extra shroud needed to be installed - I did this and migrated the cables for the 4 whitening boards from the P1 to P2 connectors. So until the new Acromag box is installed, we have no control over the whitening gains (slow channels), but do still have control over the whitening filter enable/disable (controlled by fast BIO). I am thinking about the easiest way to test the latter - I think the ambient PD dark noise level is too low to be seen above ADC noise even with the whitening enabled, and setting up drive signals to individual channels is too painful - maybe with +45dB of whitening gain, the (z,p) whitening filter shape can be seen with just PD/demod chain electroncis noise.
This morning, I wanted to move the existing cables going to the P1 connectors of the iLIGO whitening boards to the P2 connector, to test the modifications made to allow whitening stage switching. Unfortunately, I found that the shrouds werent installed. Where can I find these?
> Looking through the manual, I found a recommendation (pg10) that the "IN-" terminal of the Acromag ADC units be tied to the "RTN" pins on the same units. I don't know if this preserves the differential receiving capability of the Acromag ADCs
I suppose, we loose the differential capability of an input if the -IN is connected to whatever defined potential. We should check if the channels are still working as a true differential or not.
2. If the multi bit operation is too complicated to solve, we can use EPICS Calc channels to breakout a value to bits and send the individual bits as same as the other individual binary channels.
Any other ideas? The problem persists and it's annoying that the IMC cannot be locked.
With Chub's help, most of the problems have been resolved. Summary: I judge that we are good to go ahead with an install tomorrow.
Since we don't immediately need the CM board, I say we push ahead with the install - at least that will restore the ability to lock PRMI / DRMI. Then we can debug these issues in situ - I'm certain the issue is related to the EPICS/Modbus setup and not the hardware because I verified the physical channel map using the Acromag windows utility.
I bench tested the functionality of all the c1iscaux Acromag crate channels. Summary: we are not ready for a Monday install, much debugging remains.
I am leaving the crate powered (by bench supplies) in the office area so I have the option to work remotely on this.
I borrowed an old-looking Variac variable transformer from the power supplies cabinet along the y-arm. It is currently in the TCS lab.
We want to know that we can lock the interferometer with the ALS beat note being generated by beating IR pickoffs (rather than the vertex green transmission). The hope is also to make the ALS system good enough that we can transition the CARM offset directly to 0 after the DRMI is locked with arms held off resonance.
Attachment #1: Shows the layout. The realized MM is ~36 %. c.f. the 85% predicted by a la mode. It is difficult to optimize much more given the tight layout, and the fact that these fast lenses require the beam to be well centered on them. They are reasonably well aligned, but I don't want to futz around with the pointing into the doubling crystal. Consequently, I don't have much control over the pointing.
Attachment #2: Shows pictures of the fiber tips at both ends before/after cleaning. The tips are now much cleaner.
The BeatMouth NF1611 DC monitor reports ~580 mV with only the EY light incident on it. This corresponds to ~60 uW of light making it to the photodiode, which is only 25% of what we send in. This is commensurate with the BS loss + mating sleeve losses.
To find the beat between PSL and EY beams, I had to change the temperature control MEDM slider for the EY laser to -8355 cts (it was 225 cts). Need to check where this lies in the mode-hop scan by actually looking at the X-tal temperature on the front panel of the EY NPRO controller - we want to be at ~39.3 C on the EY X-tal, given the PSL X-tal temp of ~30.61 C. Just checked it, front panel reports 39.2C, so I think we're good.
EY enclosure was closed up and ETMY Oplev was re-enabled after my work. Some cleanup/stray beam dumping remains to be done, I will enlist Chub's help on Monday.
We scoped out the 1Y3 rack this morning to figure out what needs to be done hardware wise. We did not think about how to power the Acromag crate - the LSC rack electronics are all powered by linear supplies and not Sorensens, and the linear supplies are operating at pretty close to their maximum current-drive. The Acromag box draws ~3A of current from the 20 V supply, not sure what the current draw will be from the 15 V supply. Options:
I'm going with option #2 unless anyone has strong objections.
At ~1am PDT today, all the MC1 shadow sensor readbacks (fast CDS channels and Slow Acromag channels, latter not shown here) went to negative values. Of course a negative value makes no sense. After ~3 hours, they came back to positive values again. But since then, the shadow sensor RMS noise has been significantly higher in the >20 Hz band, and there are frequent glitches which kick the suspension. The IMC has been having trouble staying locked. I claim that this has to do with the Satellite box.
No action being taken now while I work on the ALS. In the past the problem has fixed itself.
Couple IR light into fiber with good MM at EY
I grab 2 hours of the PD measurements using dlData_simple.ipynb in the misaligned state.
I get pretty much a normally distributed reading without drifts (Attachements 1 and 2).
The error in the reading is ~ 0.5%.
I am pretty sure this amount of noise is enough to explain the big noise in the Loss figure measurement.
The reason is that the loss formula is #(1-P_Locked/P_Misaligned+T1)-T2) where T1 and T2 are the transmissions of the ITM and ETM.
The average of the ratio P_Locked/P_Misaligned is ~ 1.01 for a loss figure of ~ 100ppm.
The standard deviation of the ratio is ~ 1% which is also the standard deviation of the expression in the brackets.
The average of this experssion however is ~ 0.01.
The reduction of the mean amplifies the error in the loss measurments by a factor of a few 10s!
The requirement on the phase noise on the direct backscatter from the OMC back into the SRM is that it be less than @ 100 Hz, for a safety factor (arbitrarily chosen) of 10 (= 20dB below unsqueezed vacuum). Assuming 5 optics between the OMC and SRM which contribute incoherently for a factor of sqrt(5), and assuming a total of 1 ppm of the LO power to be backscattered, we need the suspensions to be moving @ 100 Hz. This seems possible to realize with single stage suspensions - I assume we get f^4 filtering from the pendulum at 100 Hz, and that there is an additional 80 dB attenuation (from the stack) of the assumed 1 micron/rtHz motion at 100 Hz, for an overall 160 dB attenutaiton, yielding 10^-14 m/rtHz at 100 Hz.
This is the same calculation as I had posted a couple of months ago (see elog that this is a reply to), except that Koji pointed out that the LO power is expected to dominate the (carrier) power incident on the OMC cavity(ies). So the more meaningful comparison to make is to have the x-axes of the plots denote the backscatter fraction, rather than the LO power. One subtlety is that because the phase of the scattered field is random, the displacement-noise induced phase noise could show up in the amplitude quadrature. I think that in these quadrature field amplitude units, the RIN and phase noise are directly comparable but I might have missed a factor of 2*pi. But in the worst case, if all the phase noise shows up in the amplitude quadrature, we end up being only ~10dB below unsqueezed vacuum (for 1 ppm backscatter).
For the requirement on the noise in the intensity quadrature - I think this is automatically satisfied because the RIN requirement on the incident LO field is in the mid 10^-9 1/rtHz regime.
ML2013 is unable to open Simulink on any of the workstations. We decided to make the default version of Matlab R2015b (the default of the version of RCG we are using).
I commenced the procedure of the migration, starting with making a tagged commit of the current running simulink models. A local backup was also made, plus we have the usual chiara-based backup so I think we're in good hands.
Currently the branch and tag are protected - once we verify that everything works as expected post migration, I will open it up. I changed the directory structure of the models, need to confirm that the rtcds compilers don't have any hardcoded paths which may break due to my change.
The symlink to Matlab R2013 was deleted and a new symlink to R2015b was made. I activated the license using the Caltech campus license. Now running matlab from shell starts up R2015b . Simulink even works 😲 .
cdsutils is not working on rossa.
Import cdsutils produces this error:
In : import cdsutils
OSError Traceback (most recent call last)
<ipython-input-2-949babce8459> in <module>()
----> 1 import cdsutils
/ligo/apps/linux-x86_64/cdsutils-480/lib/python2.7/site-packages/cdsutils/__init__.py in <module>()
---> 55 import awg
56 except ImportError:
/ligo/apps/linux-x86_64/cdsutils-480/lib/python2.7/site-packages/cdsutils/awg.py in <module>()
---> 32 import sys, numpy, awgbase
33 from time import sleep
34 from threading import Thread, Event, Lock
/ligo/apps/linux-x86_64/cdsutils-480/lib/python2.7/site-packages/cdsutils/awgbase.py in <module>()
17 libawg = CDLL('libawg.so')
18 libtestpoint = CDLL('libtestpoint.so')
---> 19 libSIStr = CDLL('libSIStr.so')
/ligo/apps/anaconda/lib/python2.7/ctypes/__init__.pyc in __init__(self, name, mode, handle, use_errno, use_last_error)
365 if handle is None:
--> 366 self._handle = _dlopen(self._name, mode)
368 self._handle = handle
OSError: libSIStr.so: cannot open shared object file: No such file or directory
We check for unexpected drifts in the PD reading (clipping and such). We put a pickoff mirror where the PD used to be and place the PD at the edge of the table such that the beam is focused on it (see attachment).
The arms are completley misaligned. We note the time of start of measurement to be 1249086917.
The VEA laptop asia was configured to be able to connect to too many WiFi networks - it was getting conflicted in its default position at the vertex and trying to hop between networks, for some reason trying to connect to networks that had poor signal strength. I deleted all options from the known networks except 40MARS. Now the network connection seems much more stable and reliable.
We hypothesize that the systematic error in the loss measurement can come from the fact that the requirement on the alignment of the cavity mirrors is not stringent enough.
We repeat the loss measurement with 50 measurements. This time we change the thresholds for the error signals of the dither-align in the measureArmLoss.py file from 0.5 to 0.3.
We repeat the analysis done before:
We plot the reflected power of the two states on top of each other:
This time it appears there was no drift. The histogram of the loss measurement:
The mean is 104ppm and the variation is 27%.
What I notice this time is that the PD readings in the aligned and misaligned states are anti-correlated. This is also true in the previous run (where there was drift) when looking in the short time scales. I plot several time series to demonstrate:
I wonder what can cause this behaviour.
I analyze the 100 reps loss measurement of the Y arm using the AnalyzeLossData.ipynb notebook.
The mean of the measured loss is ~ 100ppm and the variation between the repititions is ~ 27%.
In the real measurement the misaligned and locked states are repeatedly switched between each other. I plot the misaligned and locked PD readings seperately over time.
There seems to be a drift that is correlated between the two readings. This is probably a drift in the power after the MC2. To verify, I plot the ratio between those readings and find no apparent drift:
The variation in the ratio is less than 1%. The loss figure, computed to be 1 minus this ratio times a big number, give a much worse variation. I plot the histogram of the loss figure at each repitition (excluding extremely bad measurements):
The mean is ~ 100ppm. And the variation is ~ 27%.
I want to collect some data with the arms locked to investigate the possibility/usefullness of having seismic feedforward implemented for the arms (it is already known to help the IMC length and PRC angular stability at low frequencies). To facilitate diagnostics I modified the file /users/Templates/Seismic/Seismic_vs_TRXTRYandMC.xml to have the correct channel names in light of Lydia's channel name changes in 2016. Looking at the coherence data, the alignment of the cartesian coordinate system of the Seismometers at the ends and the global interferometer coordinate system can be improved.
I don't know if for the MISO filter design if there is any difference in using TRX/TRY as the target, or the arm length control signal.
Data collection started at 1249018179. I've setup a script running in a tmux shell to turn off the LSC enable in 2 hours.
We run a loss measurement on the Y arm with 50 repetitions.
I've put the analog camera back and disconnected the 151 unit GigE. But I ran out of time and wasn't able to replace the beamsplitter. I've put all the equipments back to the place where I took them from. The chopper and beam dump mount, that Koji had got me for the scatterometer, are kept outside, on the table I was working on earlier, in the control room. The camera lenses, additional GigEs, wedge beamsplitter, 1050nm LED and all related equipments are kept in the GigE box. This box was put back into CCD cameras' cabinet near the X arm.
Note: To clean stuff up, I had entered the lab around 9.30pm on Monday. This might have affected Yehonathan's loss measurement readings (until then around 57 readings had been recorded).
Sorry for the late update.
I'd like to confirm that the IR ALS scheme will work for locking. The X-arm performance so far has been encouraging. I want to repeat the characterization for the Y arm. So I inspected the layout on the EY table, and made a list of characterization tasks. The current EY beam routing is difficult to work with, and it will definitely benefit from a re-do. However, I don't know how much time I want to spend re-doing it, so for a start, I will just try and couple some amount of light into a fiber and bring it to the PSL table, and see what noise performance I get.
Attachment #1: Photo of the current beam layout. The powers indicated were measured with the Ophir power meter.
Attachment #2: A candidate mode-matching solution, given the constraints outlined above. It isn't great, with only 85% modematching even theoretically possible. The lenses required are also pretty fast lenses. But I think it's the best possible without a complete overhaul of the EY layout. I'm still waiting for the lens kit to arrive, but as soon as they get here, I will start this work.
vanna --> QIL.
gautam 20190804: The GPIB module + power supply were returned to me by Duo ~5pm today at the 40m.
4 deg is not an optimized number optimized for criteria, but to keep the cavity short width to 0.1m. But the justification of 4deg is found in Section 3 and 4 of T1000276 on Page 4.
Question for Koji: how is the aLIGO OMC angle of incidence of ~4 degrees chosen? Presumably we want it to be as small as possible to minimize astigmatism, and also, we want the geometric layout on the OMC breadboard to be easy to work with, but was there a quantitative metric? Koji points out that the backscatter is also expected to get worse with smaller angles of incidence.
Chub brought the replacement Supermicro we ordered to the 40m today. I stored it at the SW entrance to the VEA, along with the other Supermicro. At the time of writing, we have, in hand, two (unused) Supermicro machines. One is meant for EY and the other is meant for c1psl/c1iool0. DDR3 RAM and 120 GB SSD drives have also been ordered, but have not yet arrived (I think, Chub, please correct me if I'm wrong).
Update 20190802: The DDR3 RAM and 120 GB SSD drives arrived, and are stored in the FE hardware cabinet along the east arm. So at the time of writing, we have 2 sets of (Supermicro + 120GB HD + 4GB RAM).
We should ask Chub to reorder several more SuperMicro rackmount machines, SSD drives, and DRAM cards. Gautam has the list of parts from Johannes' last order.
We need to determine the geometry (= round-trip length and RoC of curved mirrors) of the OMC cavities for the 40m BHD experiment. Sticking to the aLIGO design of a 4 mirror bowite cavity with 2 flat mirrors and 2 curved mirrors, with a ~4deg angle of incidence, we need to modify the parameters for the 40m slightly on account of our different modulation frequencies. I've setup some infrastructure to do this analytically - even if we end up doing this with Finesse, it is useful to have an analytic calculation to validate against (also not sure if Finesse can calculate HOMs up to order 20 in a reasonable time, I've only seen maxtem 8).
Attachment #1: Heatmap of the OMC transmission for the following fields:
The code used for the ABCD matrix calcs have been uploaded to the BHD modeling GIT (but not the one for making this plot, yet, I need to clean it up a bit). Some design considerations have also been added to our laundry list on the 40m wiki.
One of the biggest challenges in LIGO is reducing the alignment control noise. If you haven't worked on it for at least a few years, it probably seems like a trivial problem. But all versions of LIGO since 2001 have been limited by ASC noise below ~50 Hz.
I think the 40m IMC is a good testbed for us to try a few approaches towards mitigating this noise in LIGO. The following is a list of steps to take to get there:
I think that steps 1-6 are well within our existing experience, but we should do it anyway so as to reduce the IMC beam motion at low frequencies, and also to reduce the 10-100 Hz frequency noise as seen by the rest of the interferometer.
Steps 7-8 are medium hard, but we can get some help from the CSWG in tackling it.
Step is pretty tough, but I would like to try it and also get some help from MLWG and CSWG to address it.
1. X arm is totally misaligned in order to measure the Y arm loss using the reflection method. Each measurement round consists of measuring the reflected power when the Y arm is aligned and when it is misaligned.
2. The measurement script used is /scripts/lossmap_scripts/armLoss/measureArmLoss.py. It generates a log file in the /logs folder specifying the alignment and misalignment times.
3. The data extraction script dlData.py processes the raw data in the log file and creates a hdf5 file in the /Data folder conataining the data of the measurement it self.
4. dlData.py labels the the aligned and misaligned datas incorrectly when the number of measurement is odd. I use only even number of measurements then.
5. In order to clip the chaotic transition between the aligned and misaligned states I use tDur attribute smaller than the actual sleep time used in the measurement script itself.
6. plotData.py (written by Gautam) and AnalyzeLossData.ipynb (written by me) can be used to calculate the loss and to plot some analyses (see https://nodus.ligo.caltech.edu:8081/40m/14568). They give roughly the same answer. The descripancy can be explained by the different modulation and ITM transmissions used.
7. I take a measurement of 8 repeatitions. I plot the measured reflected power alternating between the aligned and misaligned states.
I find that the reflected power is repeatable to within 1%.
This is consistent with the transmission data plotted here which is also repeatable to within 1%.
8. I take an overnight measurement of 100 repeatitions.
I've started setting up the last new rackmount SuperMicro as a dedicated server for the GigE cameras. The new machine is currently sitting on the end of the electronics test bench. It is assigned the hostname c1cam at IP 192.168.113.116 on the martian network. I've installed Debian 10, which will be officially supported until July 2024.
I've added the /cvs/cds NFS mount and plan to house all the client/server code on this network disk. Any dependencies that must be built from source will be put on the network disk as well. Any dependencies that can be gotten through the package manager, however, will be installed locally but in an automated way using a reqs file.
I brought one CPU (Dell T3500) and one 28" monitor from Mike Pedraza's office in Downs to the 40m. It is on Steve's desk right now, pending setup. The machine already has Solidworks and Altium installed on it, so we can set it up at our leisure. The login credentials are pasted on the CPU with a post-it should anyone wish to set it up.
I think rana did some more changes to this workstation to make it useful for commissioning activities - but the MEDM screens were still white blanks. The problem was that the firewalld wasn't disabled (last two steps of the KThorne setup wiki). I disabled it. Now donatella can run MEDM, ndscope and StripTool. DTT doesn't work to get online data because of a "Synchronization Error", I'm not bothering with this for now. I think Kruthi successfully demonstrated the fetching of offline data with DTT.
These spectra were taken with the arm cavity length locked to the PSL frequency using POX as an error signal, and the EX laser frequency locked to the XARM cavity length by the analog PDH servo at EX, so there is no feedback control with the ALS beat signal as an error signal.
I'll keep developing the camera server on a parallel track using the "new_..." directory naming convention. One thing I forgot to note is that the new pylon/pypylon packages require Python 3, so will not work with any of the 2.7 scripts. All of the environment I need to set up is exclusively Python 3. I won't change anything in the Python 2.7 environment in current use.
Also, I found the source of the bit resolution issue: Joe B's code loads a set of initialization parameters from a config file. One of them is "Frame Type = Mono8" which sets the dynamic range of the stream. I'll look into how this should be changed.
Since there are multiple SURF projects that rely on the cameras:
Somehow I never got around to doing the pixel sum thing for the new real data from the GigE. Since I have to do it for the presentation, I'm putting up the results here anyway. I've normalized this and computed the SNR with the true readings.
SNR = (power in true readings)/ (power in error signal between true and predicted values)
Attachment #2 is SNR of best performing CNN for comparison.
My changes were necessary because the grabHDR.py script was throwing python exceptions, whereas it was running just fine before Jon's changes. We can move the "new_*" dirs to the default once the SURFs are gone.
Let's freeze the camera software config in this state until next week.
At the lab meeting today, Rana suggested that I use the Pylon app to collect more data if that's what I need. Following this, Jon helped me out by updating the pylon version and installing additional software to record video. Now I am collecting data at
Consequently I have dithered the MC2 optic from around 9:00 PM.
I upgraded Pylon, the C/C++ API for the GigE cameras, to the latest release, 5.2.0. It is installed in the same location as before, /opt/rtcds/caltech/c1/scripts/GigE/pylon5, so environment variables do not change. The old version, 5.0.12, still exists at opt/rtcds/caltech/c1/scripts/GigE/backup_pylon5.
The package contains a GUI application (/bin/PylonViewerApp) for streaming video. The old version supports saving still images, but Milind discovered that the new version supports saving video as well. This required installing a supplementary package supporting MPEG-4 output.
Basler's GUI application is launched from the terminal using the alias pylon. I've tested it and confirm it can save both videos and still-image formats. I recommend to also try grabbing images using this program and check the bit resolution. It would be a useful diagnostic to know whether it's a bug in Joe B.'s code or something deeper in the camera settings.
Moved the unstick.py code to the ifotest repository here. It now handles signals like those generated by Ctrl-C and so forth. It can still be run as python unstick.py <machine1> <machine2> etc.
Summary: I calibrated MC2 pitch and yaw offsets to spot position in mm. Here's what I did:
Results: In the pitch/yaw vs pitch_offset/yaw_offset graph attached,
I have been trying a couple of HDR algorithms, all of them seem to give very different results. I don't know how suitable these algorithms are for our purpose, because they are more concerned with final display. I'm attaching the HDR image I got by modifying Jigyasa's code a bit (this image has been be modified further to make it suitable for displaying). Here, I'm trying compare the plots of images that look similar. The HDR image has a dynamic ratio of 700:1
PS: 300us_image.png file actually looks very similar to HDR image on my laptop (might be an issue with elog editor?). So I'm attaching its .tiff version also to avoid any confusion.
I succeeded in locking the PSL frequency to the XARM cavity length, with 9 pm RMS (Attachment #1) motion below 1 kHz, by actuating on MC2 to change the IMC length. The locks were pretty stable (~20 minutes) - the dominant cause of lockloss was the infamous ETMX drifting problem.
My main motivation here is to make some measurements and investigate the SoCal idea using a toy system, i.e. a single arm cavity controlled using ALS, so that's what Craig and I will attempt next.